### v3

# arXiv:2509.21812v3[econ.TH] 02 Mar 2026

Abstract
We investigate Ekici (2024b)’s multi-center allocation problems, focusing on fairness in this context. We introduce three fairness notions that respect centers’ priorities: internal fairness, external fairness, and procedural fairness. The first notion eliminates envy among agents within the same center, the second prohibits envy across different centers, and the third rules out envy from an ex-ante perspective through agents’ trading opportunities.
We provide two characterizations of a natural extension of the top-trading-cycles mechanism (TTC) through our fairness notions. Precisely, we show that in the presence of strategy-proofness and pair efficiency, internal fairness and external fairness together characterize TTC (Theorem 1). Additionally, strategy-proofness combined solely with procedural fairness also characterizes TTC (Theorem 2). Furthermore, by adding internal fairness, we establish our third TTC characterization, by relaxing Ekici (2024b)’s queuewise rationality to another voluntary participation condition, the center lower bound (Theorem 3). Finally, we define a core solution within this model and characterize it through TTC (Theorem 4).
Our findings offer practical insights for market designers, particularly in contexts such as international cooperation in medical programs and worker exchange programs.
JEL classification: C78; D61; D47.
Keywords: multi-center allocation problems; fairness; pair efficiency; strategy-proofness; top trading cycles.
1 Introduction
We consider allocation problems with incomplete priorities: Agents and objects are partitioned regionally, and priorities are internally induced only within each region. The considered model is introduced by Ekici (2024b), and it is a generalization of Shapley and Scarf (1974)’s housing markets model and Hylland and Zeckhauser (1979)’s house allocation model. This model covers important practical applications, e.g., organ transplant network and worker exchange problem.
In organ transplantation networks (e.g., for living donor liver or kidney transplants), collaboration across regions is becoming increasingly important (see, for instance Webpage ENCKEP, 2024; Webpage United Network for Organ Sharing, 2024a, b): While kidney and liver transplantation programs typically operate regionally, cross-program cooperation is necessary for better outcomes, such as helping more patients receive organs and improving match quality. This kind of collaboration is becoming more common (Rees et al., 2024). To address this issue, Ekici (2024b) introduces a new model called multi-center allocation problems, focusing on organ transplantation networks. [^1] A partial list of papers that study on the context of multi-center kidney exchange and dynamic matching is Roth et al. (2007); Abraham et al. (2007); Rees et al. (2017); Agarwal et al. (2019); Akbarpour et al. (2024). In this model, agents (patients) and objects (organs) are partitioned into different medical centers, and each center follows a strict order to rank its members, originally designed for running its own transplantation program locally. By accounting for the complex structure of this model, we can incentivize organ transplantation programs to operate more globally, improving efficiency while ensuring voluntary participation and fairness.
In worker exchange programs, a worker may be replaced by a new one at his home organization. A few examples include the Commonwealth Teacher Exchange Programme (Dur and Ünver, 2019), teacher reassignments in France (Combe et al., 2022), and job rotation in Japan (Yu and Zhang, 2020). In such exchanges, the key question is how to reassign workers in an efficient and fair manner. The model we consider can also be applied to this type of allocation problems, where employees’ assigned positions are switched. In this context, agents (employees) and objects (job positions) are distributed across different departments within an organization (a firm or a public sector). Each department has information only about its own employees, resulting in an internal priority that may reflect the department’s evaluation of its employees.
In the context of organ transplantation networks, Ekici (2024b) introduces a new voluntary participation condition, queuewise rationality, which states that an agent ordered in the $q$-th position in a center is guaranteed to receive an object at least as good as his $q$-th-favorite object in that center. Furthermore, he demonstrates that a natural extension of the top-trading-cycles mechanism (TTC in short) is characterized by pair efficiency, strategy-proofness, and queuewise rationality. Consequently, one can conclude that TTC is outstanding in this context, as it is efficient, strategically robust, and ensures voluntary participation.
1.1 Overview of the paper
In this paper, unlike Ekici (2024b), we primarily focus on the issue of fairness: Fairness is particularly important for implementation, as participants may reject an unfair mechanism and refuse to use it. One relevant example of this happened in school choice in New Orleans, where TTC was used in 2010 but was later replaced by the deferred acceptance mechanism, because TTC was considered unfair in the context of school choice (for details, see Abdulkadiroǧlu et al., 2020). That is to say, before implementing a mechanism, we need to verify whether it satisfies certain fairness criteria. To address it, we propose several fairness notions, derived from centers’ priorities, aimed at eliminating agents’ envies.
We begin by proposing two notions that aim to eliminate agents’ envies over allocations from an ex-post perspective. The first one, internal fairness, states that an agent should never envy another agent who is from the same center but has a lower priority than himself. Next, to determine whether envy between agents from different centers is justifiable, we introduce our second fairness property, external fairness, inspired by Morrill (2015). Following Morrill (2015), if agent $i$ envies agent $j$ from a different center, we check whether reassigning $j$’s allotment to $i$ could potentially harm any agents from $i$’s center who are ranked higher than $i$. If so, we consider $i$’s envy unjustifiable.
The third notion, procedural fairness, addresses eliminating agents’ envies over trading opportunities from an ex-ante perspective. We introduce a notable notion, agents’ trading opportunity sets: Each agent’s trading opportunity set consists of objects belonging to the agent’s center, respecting the center’s priority for that agent. Agents can use objects from their trading opportunity set for direct consumption or for the exchange with other agents. procedural fairness says that no set of agents can reach preferred objects by trading only within their trading opportunity sets.
Our main findings provide several TTC characterizations through our fairness notions. Specifically, we show that: (1) along with strategy-proofness and pair efficiency, internal fairness and external fairness characterize TTC (Theorem 1); and (2) together with strategy-proofness, procedural fairness also characterize TTC (Theorem 2). Our results further highlight the exceptional performance of TTC from a fairness perspective. Moreover, we consider an additional voluntary participation condition, the center lower bound, adapted from Chen and Feng (2025), which is weaker than Ekici (2024b)’s queuewise rationality. We also provide another TTC characterization by replacing external fairness with the center lower bound (Theorem 3). Finally, we also explore “(strict) core” solution in our context. We modify the standard blocking notion to fit our model and correspondingly have a definition of our core solution, referred to as proper core, and show that it is single-valued and coincides with TTC (Theorem 4).
Overall, together with Ekici (2024b)’s result, our results suggest that TTC perform surprisingly well across four common objectives: Fairness, efficiency, incentives, and voluntary participation.
Organization
The rest of the paper is organized as follows: In the next section, we provide an overview of the related literature. In Section 3, we introduce the basic model, the mechanisms with their properties, and a formal description of TTC. In Section 4, we first characterize TTC through our first two fairness properties (Theorem 1). We then provide another TTC characterization (Theorem 2) based on our third fairness property. In Section 5, we distinguish our results from those of Ekici (2024b) and offer an alternative TTC characterization (Theorem 3) by weakening Ekici (2024b)’s queuewise rationality. In Section 6, we discuss stability notions within our model, define a core solution in this context, and characterize it via TTC (Theorem 4). Finally, we conclude the paper in Section 7 with some final remarks. Appendix A contains all proofs omitted from the main text, and we demonstrate the independence of our properties in Appendix B.
2 Related literature
The paper closest to ours is Ekici (2024b) and we further discuss the relation between his and ours in Section 5. The main result of this paper contributes to the literature that studies TTC from an axiomatic perspective, e.g., Ma (1994). Morrill and Roth (2024) provide a comprehensive survey on this topic.
Fairness
Our paper is related to a growing literature that investigates fairness notions that are compatible with efficiency. Our first notion, internal fairness, is closely related to weak fairness (Svensson, 1994) and respect for priority (Coreno and Balbuzanov, 2022), both of which ensure the absence of justified envy. The main difference lies in implications: while their fairness properties are compatible with efficiency, they only result in dictatorial rules, which are often viewed as inequitable. In contrast, our property allows for non-dictatorial mechanisms, highlighting a key distinction.
Our second notion, external fairness, aligns with the idea that an agent with envy must propose an alternative allocation. Work in this vein includes concepts such as justness (Morrill, 2015), partial fairness (Dur et al., 2019), essential stability (Troyan et al., 2020), and legality (Ehlers and Morrill, 2020). Most papers in this literature are loosely based on the idea that priorities can be violated in certain situations, such as when agents “consent” to having their priority violated if they cannot benefit from it, while other agents can be better off through the violation of priorities. This reveals a foundational difference between those papers and ours: In their setting, priorities are complete, whereas in ours, priorities are incomplete (recall that each center’s priority is only over its own members). As a result, the key challenge in our model lies in defining fairness with incomplete priorities, rather than focusing on priority violations. Nevertheless, external fairness shares similar ideas with these studies, as it also emphasizes the consequences of agents’ envies—specifically, how allowing one agent’s envy can trigger a chain of reassignments (see Example 1). While similar in the sense that both consider more than one step ahead, external fairness fundamentally differs from notions like essential stability. Essential stability takes an optimistic approach: When anticipating future steps, they focus on the best possible outcomes, assuming agents achieve their best possible allotments. In contrast, external fairness adopts a pessimistic approach, focusing on excluding all potential harms to other agents. More broadly, external fairness contributes to the literature on defining stability in scenarios where agents anticipate more than one step of blocking. The standard concept in this literature is called farsightedness (Harsanyi, 1974): An outcome is farsightedly stable if no series of blocks exists that culminates in better outcomes for every agent participating in it. This concept is recently explored into priority-based allocation problems, e.g., Dogan and Ehlers (2023) and Atay et al. (2025).
Our concept of procedural fairness offers new insights by emphasizing potential unfairness in the allocation process rather than in the allocations themselves. Relevant literature on procedural fairness includes Klaus and Klijn (2006) and Cheng and Yu (2021). Different from ours, Klaus and Klijn (2006) focus on random allocations. They study probabilistic stable mechanisms that meet their requirement: whenever each agent has the same probability to move at a certain point in the procedure that determines the final probability distribution. Cheng and Yu (2021) study priority-based allocation problems and introduce a control system where the control rights of each object are concentrated based on priorities and no coalition has redundant control rights, and propose a non-empty core concept based on the control system. This paper differs from their work in the following aspects: First, in our setting, priorities are incomplete, whereas in theirs, objects are endowed with complete priorities. Second, in their paper, the control right of each agent relies on that agent’s received allotment, while the trading opportunity set of each agent does not. Third, unlike their system, which allows for both direct and indirect control, our opportunity set focuses solely on direct control (i.e., consuming your own or doing exchange with other agents) due to the practical challenges of implementing indirect control in kidney exchanges, especially among agents from different centers. Another related notion is individual trade stability (Pápai, 2013). This notion somewhat is between our external fairness and procedural fairness. It states that agent $i$’s envy to $j$ is not justified if $j$ obtains the object $i$ desires through a trade with a third agent, $k$, who has a higher priority than $i$ for that object. In other words, individual trade stability eliminates envies by accounting for agents’ tradings rather than depending on objects’ priorities.
Incomplete priorities
More broadly, our paper contributes to the topic of allocation problems with incomplete priorities, e.g., Dur and Ünver (2019). In this type of problems, objects only have rank over partial agents and thus their priorities may be incomplete. Given this restriction, how to define fairness is crucial. In Dur and Ünver (2019), their fairness notion, respect for internal priorities is related to objects’ eligibility quotas as in their model, objects’ capacities may vary. However, in our setting, we cannot change the number of objects due to the constraints in reality, e.g., the supply of organs are scarce and may be fixed through a period. Instead, we focus on fairness within a framework where the size of the problem is fixed.
Our paper also contributes to the growing literature focusing on allocation problems with complex ownership or property rights (Ekici, 2013; Balbuzanov and Kotowski, 2019; Sun et al., 2020; Zhang, 2020; Cheng et al., 2023). In these types of problems, ownership (or agents’ rights) can often be translated into partial priorities: owners have strictly higher priority than other agents, while non-owners are treated as indifferent among themselves. This creates a challenge similar to the one we address: It is hard to compare two agents when there is no strict order to rank them. In other words, all face the same question: How to expand priorities into linear orders in a fair manner. We further discuss this point in Section 6.
3 Model
3.1 Multi-center allocation problems
The basic model setup is same to Ekici (2024b). There is a finite set of centers, $C$. Each center $c\in C$ consists of a finite set of agents (e.g., patients or workers) $N_{c}$ and a finite set of heterogeneous objects (e.g., organs or job positions) $O_{c}$, with $|N_{c}|=|O_{c}|$. [^2] In this paper, we do not consider the outside option because we want to keep balance in our model, we further discuss this point in Section 7. Each center $c$ has a strict priority $\succ_{c}$ over $N_{c}$. Each center’s priority of its members reflects some form of justice—whether based on need, queue time, or other criteria. However, the design of such priorities is beyond the scope of this paper, and we therefore assume that the priority structure is exogenously given.
Let $N\equiv\cup_{c\in C}N_{c}$ and $O\equiv\cup_{c\in C}O_{c}$ denote the set of all agents and all objects, respectively. Let $n\equiv|N|=|O|\geq 2$. For each agent $i\in N$, let $c(i)$ represent the center to which they belong to, i.e., $i\in N_{c(i)}$. An allocation is a bijection $x:N\to O$. Let $X$ denote the set of all allocations. For each $x\in X$ and each $i\in N$, $x_{i}\equiv x(i)$ denotes agent $i$’s allotment at $x$. Let $R=(R_{i})_{i\in N}$ be a preference profile over $O$, where $R_{i}$ denotes the strict preference of agent $i$. The strict part associated with $R_{i}$ is denoted by $P_{i}$. [^3] That is, for all $a,b\in O$, $a\mathbin{R_{i}}b$ if and only if $a\mathbin{P_{i}}b$ or $a=b$. Let $\mathcal{R}$ be the set of all strict preferences. We use the standard notation $(R^{\prime}_{i},R_{-i})$ to denote the profile obtained from $R$ by replacing agent $i$’s preference relation $R_{i}$ with $R^{\prime}_{i}\in\mathcal{R}$.
A multi-center allocation problem (or a problem in short) is formed by $((N_{c},O_{c},\succ_{c})_{c\in C},R)$. As the $(N_{c},O_{c},\succ_{c})_{c\in C}$ remain fixed throughout, we will simply denote the problem by $R$. Thus, the strict preference profile domain $\mathcal{R}^{N}$ also denotes the set of all problems.
Note that our model extends various classical object allocation models: [^4] Note that if $n=2$, then our model reduces to either the Shapley-Scarf housing market model or the house allocation model.
  • If for each $c\in C$, $|N_{c}|=|O_{c}|=1$, then our model equals the Shapley-Scarf housing market model (Shapley and Scarf, 1974); and
  • if $|C|=1$, then our model is the modified house allocation model. Note that in Hylland and Zeckhauser (1979)’s original model, there is no priority over agents.
3.2 Mechanisms and their properties
A mechanism is a function $f:\mathcal{R}^{N}\to X$ that associates with each problem $R$ an allocation $f(R)$. For each $i\in N$, $f_{i}(R)$ denotes agent $i$’s allotment at $f(R)$.
Here, we introduce several properties of mechanisms that are related to fairness.
The first one is strategy-proofness, which is most frequently used in the literature on mechanism design.
Strategy-proofness. For each $R\in\mathcal{R}^{N}$, each $i\in N$, and each $R^{\prime}_{i}\in\mathcal{R}$, $f_{i}(R)\mathbin{R_{i}}f_{i}(R^{\prime}_{i},R_{-i})$.
Typically, strategy-proofness is regarded as an incentive property that eliminates an agent’s incentive to misreport his preferences individually. As a result, it is also viewed as a fairness notion, as it “ ‘leveling the playing field’ idea only indicates that sophisticated students lose their strategic rents under the new mechanism” (Pathak, 2011).
Given $R$, we say agent $i$ envies another agent $j$ at $x$ if $x_{j}\mathbin{P_{i}}x_{i}$. An ideal allocation should eliminate all agents’ envies, resulting in what is known as an envy-free allocation. However, since our model only allows deterministic allocations, such ideal situations does not exist in general. [^5] For instance, if $R_{i}=R_{j}$, then for any allocation $x$, it must be the case that either $i$ envies $j$, or $j$ envies $i$. Therefore, we consider some weaker properties. The first one is pair efficiency (Ekici, 2024a), which states that no pair of agents should envy each other.
Pair efficiency. For each $R\in\mathcal{R}^{N}$, there is no pair of agents $i,j\in N$ such that $f_{i}(R)\mathbin{P_{j}}f_{j}(R)$ and $f_{j}(R)\mathbin{P_{i}}f_{i}(R)$.
Originally, pair efficiency is viewed as an efficiency notion, which is weaker than Pareto efficiency. [^6] An allocation $x$ is Pareto efficient if there is no distinct allocation $y\neq x$ such that for each $i\in N$, $y_{i}\mathbin{R_{i}}x_{i}$. Here, one may reinterpret it as a fairness notion as it rules out the situation where two agents envy each other, aligning with the idea that swapping two objects eliminates envy (Echenique et al., 2025).
The next property reflects “a minimal level of mindfulness toward the property rights of individual agents” (Ekici, 2024b).
For each $c\in C$, and $i\in N_{c}$, let $r_{i}=1+|\{j\in N_{c(i)}\mid j\succ_{c}i\}|$ be his rank according to $\succ_{c}$, i.e., if $i$ has the $k$-highest priority at $\succ_{c}$, then $r_{i}=k$. Moreover, for each $R\in\mathcal{R}^{N}$, let $o_{i}(R)\in O_{c}$ be his $r_{i}$-preferred object in $O_{c}$, according to $R_{i}$. Formally, $o_{i}(R)$ is such that $\{o\in O_{c}\mid o\mathbin{R_{i}}o_{i}(R)\}=r_{i}$.
Queuewise rationality. For each $R\in\mathcal{R}^{N}$, each $c\in C$, and each $i\in N_{C}$, $f_{i}(R)\mathbin{R_{i}}o_{i}(R)$.
A tale of three fairness properties
In our model, since we also incorporate (incomplete) priorities over agents, along with the idea of pairwise stability in matching theory, it is natural to consider envy-freeness with respect to these priorities. However, as we mentioned earlier, the priorities in our model are incomplete: From each center, we have one priority only over its own members. As a result, determining whether an agent’s envy is justified (based on these priorities) becomes challenging due to this incompleteness. [^7] In Section 6.2, we further discuss the classical pairwise stability in our model.
First, we start with an obvious case: When two agents are from the same center. In this case, the assignment may violate the envious agent’s priority since he has a higher priority than the agent who received the object that he desired.
Given $f(R)$, agent $i$ internally justifiedly envies another agent $j$ at $f(R)$ if
  • (1)
    $i$ envies $j$ at $f(R)$;
  • (2)
    $i$ and $j$ belong to the same center, i.e., $c(i)=c(j)=c$; and
  • (3)
    $i\succ_{c}j$.
Internal fairness. For each $R\in\mathcal{R}^{N}$, there is no pair of agents $i,j\in N$ such that $i$ internally justifiedly envies $j$ at $f(R)$.
Internal fairness only eliminates envy that is not “justified” by a center’s priority of its members. This type of justified fairness, driven by priorities, arises naturally in many applications. For instance, priorities guide the allocation of medical resources to individuals based on urgency, determine eligibility for public housing based on queueing, and enable schools to rank students based on exam scores and proximity. Moreover, internal fairness captures incentive issues. [^8] In the context of school choice with walk-zone rights, Klaus and Klijn (2024) study the similar incentive issues. They refer to it as minimal-access monotonicity: No student has an incentive to give up his additional walk-zone rights. Consider the case when priorities are induced by a reserve system. If an agent with higher priority receives a less preferred object than an agent with lower priority, then the higher-priority agent may benefit by manipulating his priority, e.g., by reneging / jockeying on it, to obtain a better object. Such strategic actions commonly exist in various reserve systems and have been extensively studied in the fields of queuing theory and operations research (see, e.g., Kullback (1997) and Gross et al. (2011)). Internal fairness prohibits such strategic actions, as all agents prefer higher priorities.
Remark 1.
Weak internal fairness
One might argue that the case above is not fully justified, since $i$ may not have a claim to receive $j$’s allotment if it is outside their common center $c$. Thus, we also consider a weak version of internal fairness, which additionally requires that $j$’s allotment is also belongs to center $c$.
Weak internal fairness. For each $R\in\mathcal{R}^{N}$, there is no pair of agents $i,j\in N$ such that (1) $f_{j}(R)\in O_{c(i)}$, and (2) $i$ internally justifiedly envies $j$ at $f(R)$.
Alternatively, one can restate it with the following: For each $c\in C$, and $o\in O_{c}$, let $\succ_{o}=\succ_{c}$. We say agent $i$ has higher priority than $j$ at an object $o$ if $i\succ_{o}j$. Then, weak internal fairness says that there is no pair of agents $i,j\in N$ such that $i$ envies $j$ at $f(R)$, and $i$ has higher priority than $j$ at $f_{j}(R)$. $\diamond$
By internal fairness, we rule out envy among agents within the same center based on that center’s priority. However, envy may still arise between agents from different centers, where no priority can be directly applied to determine whether such envy is “justified.” Nonetheless, as pointed out by Cheng et al. (2023), there may be an indirect relationship derived from priorities that can be applied to eliminate such envy. To address this, we introduce our second fairness property, external fairness, which is a modification of Morrill’s justness (Morrill, 2015). The intuitive interpretation of external fairness is as follows. If $i$ envies another agent $j$ from a different center and raises a complaint about the selected allocation, we must determine whether the complaint is justifiable. [^9] Note that here we use the term “complaint,” which differs from the term “violation of priority” we used earlier to define internal fairness. The fundamental distinction between the two arises from the difficulty we face: in the case of a priority violation, we can directly use a center’s priority to justify an agent’s envy. In contrast, for a complaint, no such priority can be directly applied. Thus, violations of priorities do not exist, making it harder to determine whether the envy is justified. Since $i$ and $j$ belong to different centers, no exogenously given priority can be directly applied to judge $i$’s complaint. However, as Morrill pointed out, we can consider a hypothetical scenario: If we accept this complaint and reassign $j$’s allotment, what would be the consequence? Specifically, would another agent $k$, who belongs to the same center as $i$ and has a higher priority, be directly harmed by this reassignment? [^10] Here, by “directly harmed” we mean that $k$’s allotment is directly reassigned to $j$, which aligns with the idea of Morrill (2015). In other words, we formalize harm by considering only immediate effects. More complex chains, in which $j$’s deviation indirectly worsens $k$’s assignment without $j$ ever receiving $k$’s allotment, are not considered, as the current notion is sufficient to formalize our fairness concept and guarantee positive results. If so, agent $k$, with a higher priority, may veto $i$’s complaint. Therefore, we should reject $i$’s complaint, as we prioritize $k$’s welfare over $i$’s, according to the priority of the center they both belong to.
This idea can also be interpreted in a manner similar to the concept of farsighted stability (Harsanyi, 1974), which considers the final allocation that may result from admitting certain envies. It is also applied in matching theory to investigate weaker definitions of stability, e.g., legality (Ehlers and Morrill, 2020) and essential stability (Troyan et al., 2020).
We now present an example to highlight the key insight behind our idea.
Example 1.
There are two centers $c$ and $c^{\prime}$, along with three agents and three objects as follows. $N_{c}=\{1,2\},O_{c}=\{a,b\}$, $N_{c^{\prime}}=\{3\}$, and $O_{c^{\prime}}=\{d\}$. Agents’ preferences and centers’ priorities are given as below (agents’ “endowed objects” are underlined for emphasis).
$R_{1}:\underline{a},\underline{b},d;$
$R_{2}:d,\underline{a},\underline{b};$
$R_{3}:a,\underline{d},b;$
$\succ_{c}:2,1;\text{ and }\succ_{c^{\prime}}:3.$
Allocation 11 22 33
$x$ $b$ $d$ $a$
$y$ $b$ $a$ $d$
Consider the allocation $x$. We see that only agent $1$ envies another agent: $x_{3}=a\mathbin{P_{1}}b=x_{1}$. Agent $1$ may argue that his envy is reasonable since $a$ belongs to center $c$, the center to which agent $1$ belongs to (i.e., $1\in N_{c}$ and $b\in O_{c}$).
If we admit this envy, then agent $3$ cannot receive $a$. Consequently, agent $3$ must receive $d$; otherwise, agent $3$ would envy whoever receives $d$, and by the same logic used to justify agent $1$’s envy, we must also admit agent $3$’s envy, as $d$ belongs to $c^{\prime}$, the center to which agent $3$ belongs to. As a result, agent $2$ cannot receive $d$. In other words, agent $2$’s allotment at $x$ “depends” on agent $3$’s reported preference, in the sense of Morrill (2015). Therefore, $\{a,b\}$ must be assigned to agents $\{1,2\}$. Since agent $2\succ_{c}1$, it is reasonable to assign $a$ to agent $2$.
In summary, if we admit agent $1$’s envy at $x$, the resulting final allocation is $y$. So, agent $1$ does not benefit from his envy as $x_{1}=y_{1}=b$, and everyone else ends up worse off. This means that $1$’s envy is “vacuous” in the sense of Troyan et al. (2020), and suggests that admitting $1$’s envy at $x$ is “farsightedly unstable.” $\diamond$
To formalize our idea, we introduce Morrill’s notion of dependence, which captures potential harm arising from reassignments. An agent $i$ depends on another agent $j$ at $f(R)$ if there exists $R^{\prime}_{j}$ such that $f_{j}(R^{\prime}_{j},R_{-j})=f_{i}(R)$.
Given $f(R)$, $i$ externally justifiedly envies $j$ at $f(R)$ if
  • (1)
    $i$ envies $j$ at $f(R)$;
  • (2)
    $i$ and $f_{j}(R)$ belong to the same center, while $j$ does not, i.e., $c(i)\neq c(j)$ and $f_{j}(R)\in O_{c(i)}$; and
  • (3)
    for each agent $\in\{\ell\in N_{c(i)}\mid\ell\succ_{c(i)}i\}, kk$ does not depend on $j$ at f(R)f(R). [^11] If $\{\ell\in N_{c(i)}\mid\ell\succ_{c(i)}i\}$ is empty, then this condition is vacuously satisfied.
External fairness. For each $R\in\mathcal{R}^{N}$, there is no pair of agents $i,j\in N$ such that $i$ externally justifiedly envies $j$ at $f(R)$.
Remark 2.
Strong external fairness
Unlike internal fairness and weak internal fairness (see Remark 1), here we cannot drop the requirement that $i$ and $f_{j}(R)$ belong to the same center. Otherwise, any cross-center envy that a highest-priority agent $i$ has toward any $j(\not\in N_{c(i)})$ would be classified as externally justified (see Footnote 11), leading to an impossibility result. In other words, no mechanism can satisfy this stronger version of external fairness. $\diamond$
By external fairness, we first check whether reassigning $j$’s allotment could potentially harm any agents who have higher priority than agent $i$. If so, we treat $i$’s complaint unjustifiable and reject it. In other words, external fairness serves to rule out envies that could potentially harm another agent with higher priority. This aligns well with common sense of fairness, similar to how people generally accept that those with higher priority on a waiting list or in a queue should not be displaced.
Finally, we discuss our last fairness property, which takes an ex-ante perspective. That is, it concerns the fairness of the process itself, based on agents’ “trading opportunities.”
Agents’ trading opportunities can be understood as follows. Any allocation $x$ can be decomposed into several trading cycles (these cycles are not necessarily unique). [^12] Loosely speaking, a trading cycle is an ordered set of agents and objects that describes who trades with whom, and for what. Thus, $x$ can be viewed as being obtained through the following process: Each agent $i$ uses an object $o\in O_{c(i)}$ to trade with other agents for $x_{i}$ (if $x_{i}\in O_{c(i)}$, then $i$ effectively trades with himself). In this sense, each agent $i$ utilizes his trading opportunities to obtain $x_{i}$.
The next question is where these trading opportunities come from and how we should evaluate them to ensure fairness. Loosely speaking, the trading opportunities available to an agent may come from resources that he controls and that he can consume/utilize directly. Ideally, agents’ trading opportunities should be equal when they are treated equally ex-ante. However, this is not the case in our model: Here, agents within the same center are already treated unequally due to that center’s priority. Thus, in our context, a just distribution of agents’ trading opportunities must also respect centers’ priorities. Specifically, if two agents come from the same center, the agent with higher priority should have more opportunities (or, a weakly larger opportunity set) than the agent with lower priority. To formalize this idea, we introduce the concept of agents’ trading opportunity sets by adopting the algorithm of Pápai (2000). In the algorithm below, at the beginning, each agent with the highest priority “controls” all objects that belong to his center, meaning he has the right to use any object within that set. Then, we let each of them points to his most preferred object, and let each object points to the highest priority agent in its center. Once these agents and their corresponding trading objects are identified, the same process is applied to the remaining agents and objects. Specifically, each agent with the highest priority among the remaining agents “controls” all remaining objects that belong to his center. This iterative procedure continues, determining each agent’s trading opportunities given the preference profile.
Construction of trading opportunity sets($\mathcal{T}$)
Step 0
: Let $N^{1}=N$, and $O^{1}=O$.
Step $\tau$ $(\geq 1)$:
For each $c\in C$ with $N_{c}^{\tau}\neq\emptyset$, let $i_{c}^{\tau}\equiv\max_{\succ_{c}}N_{c}^{\tau}$ be the agent with the highest priority among the remaining agents in $N_{c}^{\tau}$. For each $c\in C$, let $i_{c}^{\tau}$ point to his best preferred object among remaining objects according to $R_{i_{c}^{\tau}}$; and let each remaining object in $O_{c}^{\tau}$ point to $i_{c}^{\tau}$. Define the trading opportunity set $T_{i_{c}^{\tau}}(R)$ as the set of objects pointing to $i_{c}^{\tau}$. Also, given agents and objects’ pointing graph, there exists at least one cycle. Remove all agents and objects involved in cycles. Let $N^{\tau+1},O^{\tau+1}$ be the remaining agents and objects, respectively. If $N^{\tau+1}$ is not empty, then proceed to step $\tau+1$.
Termination:
The algorithm terminates when no agent remains.
We use $\mathcal{T}$ to represent the algorithm above. The algorithm naturally defines each agent’s trading opportunity set because it captures a practical implementation as follows: Agents are asked, in the order of their priorities, which available object they would use for consumption or trade, to best meet their preferences. For each agent with the highest priority, his trading opportunity set equals the set of objects in their center. That is, the opportunity set for each top-ranked agent does not depend on any agent’s preference. Moreover, for each remaining agent $j$, the trading opportunity set $T_{j}(R)$ does not depend on $j$’s preference $R_{j}$, but on the preferences of agents who have higher priority than $j$.
Now, we present our idea of fairness in the process. Since the selected allocation is produced by agents either consuming or trading objects from their trading opportunity sets, unsatisfied agents have incentives to deviate from the selected allocation if they can consume or trade objects from their trading opportunity sets to achieve higher welfare. We consider these two possibilities.
Let $x\in X$ be the selected allocation. First, direct consumption. Suppose that an agent $i$ is unsatisfied at $x$ because he envies another agent $j$’s allotment $x_{j}$. If $x_{j}\in T_{i}(R)$, then it means that $i$ can directly consume $x_{j}$ from his trading opportunity set. Second, trading. Suppose that a set of agents $S$ who from different centers realize they can improve their welfare by trading objects from their respective trading opportunity sets. In this case, all agents in $S$ can deviate from $x$, by using the objects from their trading opportunity sets and proceeding with the exchange. When cooperation across multiple centers is allowed, such trading represents the most easily achievable form of exchange.
Formally, an allocation $x$ is unfairly produced at $R$ if there exist a list of centers $c_{1},\ldots,c_{L}\in C$, a list of agents $(i_{1},\ldots,i_{L})\in\Pi_{\ell=1}^{L}N_{c_{\ell}}$, and a list of objects $(o_{1},\ldots,o_{L})\in\Pi_{\ell=1}^{L}T_{i_{\ell}}(R)$ such that for each $i_{\ell}\in\{i_{1},\ldots,i_{L}\}$, $o_{i_{\ell}+1}\mathbin{P_{i_{\ell}}}x_{i_{\ell}}$ (mod $L$). Otherwise $x$ is fairly produced at $R$. Note that if there is only one center (so $L=1$), then the condition represents the direct consumption, otherwise it represents the trading among agents $i_{1},\ldots,i_{L}$.
Procedural fairness. For each $R\in\mathcal{R}^{N}$, $f(R)$ is fairly produced at $R$.
Procedural fairness offers a fresh perspective for evaluating mechanisms. For instance, while Pápai (2000) hierarchical exchange mechanisms satisfy procedural fairness, Pycia and Ünver (2017)’s trading cycles mechanisms do not. [^13] In Pycia and Ünver (2017)’s trading cycles mechanisms, brokers cannot receive the objects they control, which results in a violation of procedural fairness. This suggests the value of procedural fairness in distinguishing mechanisms. Furthermore, since trading cycles mechanisms are Pareto efficient, its violation of procedural fairness also implies that procedural fairness is logically independent of Pareto efficiency (and pair efficiency). [^14] See Appendix B for independence examples.
Note that procedural fairness implies queuewise rationality, but not vice versa.
Lemma 1.
All procedurally fair mechanisms satisfy queuewise rationality; however, the opposite is not true.
Remark 3.
Relation to other properties
If for each $c\in C$, $|N_{c}|=|O_{c}|=1$, i.e., reduced to the Shapley-Scarf housing market model, then external fairness coincides with individual rationality, and procedural fairness coincides with weak core stability; and if $|C|=1$, i.e., reduced to the house allocation model, then internal fairness, procedural fairness, and Svensson (1994)’s weak fairness all coincide with one another. [^15] Weak fairness states that an agent should never envy another agent who has a lower priority than himself. Also, in Appendix B, we provide several examples to show that our three fairness properties are logically independent for general multi-center allocation problems. $\diamond$
3.3 Top Trading Cycles
For each problem $R$, the top trading cycles mechanism (TTC) selects the allocation determined by the following top trading cycles algorithm at $R$, which we call $TTC(R)$.
Top Trading Cycles Algorithm:
Step 0
: For each $c\in C$ and each $o\in O_{c}$, let $\succ_{o}=\succ_{c}$ be the associated priority of $o$. Let $\succ_{O}=(\succ_{o})_{o\in O}$ be objects’ priority profile.
Step $\tau$ $(\geq 1)$:
Each remaining agent points to his most-preferred remaining object given $R$. Each remaining object points to the agent with the highest priority among the remaining agents, according to $\succ_{O}$. [^16] Note that, since for each center $c$ the balance constraint $|N_{c}|=|O_{c}|$ holds, if an object in $O_{c}$ remains at step $t$, then some agent in $N_{C}$ must also remain at step $t$. As a result, each object in $O_{c}$ points only to agents in $N_{c}$. There exists at least one cycle. Execute all cycles by assigning each agent involved in a cycle the object to which he points. Remove all agents and objects involved in a cycle. If some objects remain, then proceed to step $\tau+1$.
Termination:
The algorithm terminates (in at most $n$ steps) when no agent / object remains.
TTC satisfies all of our fairness properties.
Proposition 1.
TTC satisfies internal fairness (and hence weak internal fairness), external fairness, and procedural fairness.
4 Main characterizations
4.1 Internal fairness and external fairness
We characterize TTC through the first four properties that we presented previously.
Theorem 1.
A mechanism $f$ satisfies strategy-proofness, pair efficiency, internal fairness, and external fairness, if and only if $f=TTC$.
Note that from the proof of Theorem 1 below, one can see that our characterization in Theorem 1 also holds if internal fairness is weakened to weak internal fairness (see Remark 1).
4.1.1 Proof of Theorem 1
By Pycia and Ünver (2017) and Proposition 1, we know that TTC satisfies the four properties.
Thus, we only need to prove the uniqueness part. Let $f$ be strategy-proof, pair efficient, internally fair, and externally fair. The proof below is by contradiction. In short, assuming $f\neq TTC$, we use Ekici (2024a)’s notion of similarity to select a minimal preference profile on which the two mechanisms differ, and then construct a sequence of preference manipulations by agents in a non-executed cycle to derive a contradiction with efficiency of $f$.
[…]
For each pair of similarity $(x_{1},x_{2}),(y_{1},y_{2})\in\mathbb{N}^{2}$, we say $(x_{1},x_{2})<(y_{1},y_{2})$ if $x_{1}<y_{1}$ or $x_{1}=y_{1}$ and $x_{2}<y_{2}$]; and $(x_{1},x_{2})\leq(y_{1},y_{2})$ if $(x_{1},x_{2})<(y_{1},y_{2})$ or $(x_{1},x_{2})=(y_{1},y_{2})$. We also define >> and $\geq$ similarly.
[…]
Thus, $f(R)$ executes all cycles in $\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R)$, but it does not execute some cycle in $\mathcal{C}_{t_{1}}(R)$. Let $\mathcal{C}\in\mathcal{C}_{t_{1}}(R)$ be the cycle that is not executed with $|S_{\mathcal{C}}|=t_{2}$. Let $O^{t_{1}}$ be the set of objects that are remaining at step $t_{1}$ under $TTC(R)$. By the definition of TTC, we know that for each $i\in S_{C}$, $x_{i}$ is $i$’s most preferred object in $O^{t_{1}}$, i.e., for each $o\in O^{t_{1}}$, $x_{i}\mathbin{R_{i}}o$.
Claim 1.
$|S_{\mathcal{C}}|=t_{2}\geq 2$.
Proof of Claim 1.
Suppose that $|S_{\mathcal{C}}|=1$, hence it is without loss of generality to assume that $S_{\mathcal{C}}=\{i\}$ and $\mathcal{C}=(i\to x_{i}\to i)$. It implies that $x_{i}\in O_{c(i)}$. Since $f(R)$ does not execute $\mathcal{C}$, there is an agent $j\neq i$ such that $f_{j}(R)=x_{i}$. Hence $f_{j}(R)\neq x_{j}$, this implies that $j$ does not receive $x_{j}$ before step $t_{1}$ under $TTC(R)$. Since $f_{j}(R)$ is $i$’s most preferred object in $O^{t_{1}}$, $f_{j}(R)\mathbin{P_{i}}f_{i}(R)$, i.e., $i$ envies $j$ at $f(R)$. There are two cases.
Case One: $c(j)=c(i)$. In other words, $j\in N_{c(i)}$. By the selection of $R$, $f(R)$ executes all cycles in $\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R)$. By the definition of TTC, each $k\in N_{c(i)}$ with higher priority than $i$ receives $x_{k}$ from some cycle in $\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R)$. Thus, for each agent $k\in N_{c(i)}$ with $k\mathbin{\succ_{c(i)}}i$, $f_{k}(R)=x_{k}$. Therefore, $i\mathbin{\succ_{c(i)}}j$. So, (weak) internal fairness is violated.
Case Two: $c(j)\neq c(i)$, and hence $j\not\in N_{c(i)}$. By the selection of $R$, for each $R^{\prime}_{j}\in\mathcal{R}$, $\rho(R^{\prime}_{j},R_{-j})\geq t=\rho(R)$. Thus, $f(R^{\prime}_{j},R_{-j})$ executes all cycles in $\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R^{\prime}_{j},R_{-j})$. Recall that $j$ does not receive $x_{j}$ before step $t_{1}$ under $TTC(R)$. Thus, by the definition of TTC, $\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R^{\prime}_{j},R_{-j})=\cup_{\tau=1}^{t_{1}-1}\mathcal{C}_{\tau}(R)$. Hence, for each agent $k\in N_{c(i)}$ with $k\mathbin{\succ_{c(i)}}i$, $f_{k}(R^{\prime}_{j},R_{-j})=x_{k}$. Thus, $k$ does not depend on $j$. Also recall that $i\in N_{c(i)}$, $f_{j}(R)=x_{i}\in O_{c(i)}$, and $j\not\in N_{c(i)}$. Thus, $i$ externally justified envies $j$ at $f(R)$, contradicts external fairness of $f$. ∎
Note that our two fairness notions are only used to prove Claim 1.
By Claim 1, $|S_{\mathcal{C}}|=t_{2}=K\geq 2$. Without loss of generality, let $\mathcal{C}=(i_{1}\to x_{i_{1}}\to\ldots\to i_{K}\to x_{i_{K}}\to i_{1})$, and assume that $f_{i_{1}}(R)\neq x_{i_{1}}$. Let $c\equiv c(i_{1})$. That is, $i_{1}\in N_{c}$ and hence $x_{i_{K}}\in O_{c}$. Also, since $K\geq 2$, $x_{i_{1}}\not\in O_{c}$.
Let $R^{\prime}_{i_{1}}:x_{i_{1}},x_{i_{K}},\ldots$ By strategy-proofness of $f$, $f_{i_{1}}(R^{\prime}_{i_{1}},R_{-i_{1}})\neq x_{i_{1}}$; otherwise agent $i_{1}$ has an incentive to misreport $R^{\prime}_{i_{1}}$ at $f(R)$. Note that if $i_{1}$ change his report from $R^{\prime}_{i_{1}}$ to $R^{\prime\prime}_{i_{1}}$ by shifting $x_{i_{K}}$ to the top, then we have a cycle $(i_{1}\to x_{i_{K}}\to i_{1})$ in $\cup^{t_{1}}_{\tau=1}\mathcal{C}_{\tau}(R^{\prime\prime}_{i_{1}},R_{-i})$. This is because that the last higher-priority agent in $c(i_{1})$ is removed before step $t_{1}-1$ under $TTC(R^{\prime\prime}_{i_{1}},R_{-i})$, so $(i_{1}\to x_{i_{K}}\to i_{1})$ appears no later than step $t_{1}$. Then, by applying similar arguments we used to prove Claim 1, we see that this cycle is executed at $f(R^{\prime\prime}_{i_{1}},R_{-i_{1}})$, i.e., $f_{i_{1}}(R^{\prime\prime}_{i_{1}},R_{-i_{1}})=x_{i_{K}}$. Thus, by strategy-proofness of $f$, we conclude that $f_{i_{1}}(R^{\prime}_{i_{1}},R_{-i_{1}})=x_{i_{K}}$; otherwise $i_{1}$ has an incentive to misreport $R^{\prime\prime}_{i_{1}}$ at $f(R^{\prime}_{i_{1}},R_{-i_{1}})$. Therefore, $f_{i_{K}}(R^{\prime}_{i_{1}},R_{-i_{1}})\neq x_{i_{K}}$.
Let $R^{\prime}_{i_{K}}:x_{i_{K}},x_{i_{1}},\ldots$ and $R^{\prime}\equiv(R^{\prime}_{i_{1}},R^{\prime}_{i_{K}},R_{N\setminus\{i_{1},i_{K}\}})$. Since $f$ is strategy-proof, $f_{i_{K}}(R^{\prime}_{i_{1}},R_{-i_{1}})\neq x_{i_{K}}$ implies that $f_{i_{K}}(R^{\prime})\neq x_{i_{K}}$.
By dropping $x_{i_{K}}$ from top to the bottom, we obtain $R^{\prime\prime}_{i_{K}}:x_{i_{1}},\ldots x_{i_{K}}$ such that for any two objects $o,o^{\prime}\in O\setminus\{x_{i_{K}}\}$, $o\mathbin{R^{\prime\prime}_{i_{K}}}o^{\prime}$ if and only if $o\mathbin{R^{\prime}_{i_{K}}}o^{\prime}$. Thus, by strategy-proofness of $f$, $f_{i_{K}}(R^{\prime})=f_{i_{K}}(R^{\prime\prime}_{i_{K}},R^{\prime}_{-i_{K}})$.
By the definition of TTC, there is one cycle $(i_{K}\to x_{i_{1}}\to\ldots\to i_{K})=\mathcal{C}^{\prime}\in\cup_{\tau=1}^{t_{1}}\mathcal{C}_{\tau}(R^{\prime\prime}_{i_{K}},R^{\prime}_{-i_{K}})$ such that $S_{\mathcal{C}^{\prime}}=S_{\mathcal{C}}\setminus\{i_{1}\}$ and hence $|S_{\mathcal{C}^{\prime}}|=t_{2}-1=K-1$, See the figure below for the graphical explanation.
$i_{1}$$x_{i_{1}}$$i_{K}$$x_{i_{K-1}}$$x_{i_{K}}$$i_{2}$$R_{i_{1}}\penalty 10000\ /\penalty 10000\ R^{\prime}_{i_{1}}$$R^{\prime\prime}_{i_{K}}$$R_{i_{K}}\penalty 10000\ /\penalty 10000\ R^{\prime}_{i_{K}}$$R_{i_{2}}$
By the selection of $R$, $\rho(R^{\prime\prime}_{i_{K}},R^{\prime}_{-i_{K}})\geq(t_{1},t_{2})=\rho(R)$. Thus, $f(R^{\prime\prime}_{i_{K}},R^{\prime}_{-i_{K}})$ executes $\mathcal{C}^{\prime}$. Hence, $f_{i_{K}}(R^{\prime\prime}_{i_{K}},R^{\prime}_{-i_{K}})=x_{i_{1}}=f_{i_{K}}(R^{\prime})$. Since $f_{i_{K}}(R^{\prime})=x_{i_{1}}$, $f_{i_{1}}(R^{\prime})\neq x_{i_{1}}$. Again, by applying similar arguments we used above, we conclude that $f_{i_{1}}(R^{\prime})=x_{i_{K}}$. Given $f_{i_{K}}(R^{\prime})=x_{i_{1}}$ and $f_{i_{1}}(R^{\prime})=x_{i_{K}}$, we see that $f_{i_{K}}(R^{\prime})\mathbin{R^{\prime}_{i_{1}}}f_{i_{1}}(R^{\prime})$ and $f_{i_{1}}(R^{\prime})\mathbin{R^{\prime}_{i_{K}}}f_{i_{K}}(R^{\prime})$, contradicts pair efficiency of $f$. This desired contradiction completes the proof of Theorem 1.
4.2 Procedural fairness
Next, we characterize TTC through procedural fairness.
Theorem 2.
A mechanism $f$ satisfies strategy-proofness and procedural fairness if and only if $f=TTC$.
4.2.1 Proof of Theorem 2
By Pycia and Ünver (2017) and Proposition 1, we know that TTC satisfies the two properties. Thus, we only need to prove the uniqueness part. Here, we apply Sethuraman (2016)’s approach. Let $f$ be strategy-proof and procedurally fair.
We first introduce some notation. For each problem $R\in\mathcal{R}^{N}$, recall that $\mathcal{C}(R)=\cup_{\tau}\mathcal{C}_{\tau}(R)$ is the set of all trading cycles that are obtained via $TTC(R)$. For each $\mathcal{C}\in\mathcal{C}(R)$, and each involved agent $i\in S_{\mathcal{C}}$, let $e_{i}(R)$ be the object that points to agent $i$ in the cycle $\mathcal{C}$, e.g., if $\mathcal{C}=(o\to i\to o^{\prime}\to j\to o)$, then $e_{i}(R)=o$ and $e_{j}(R)=o^{\prime}$.
Next, we inductively show that for each agent $i\in N$, $e_{i}(R)$ is in his trading opportunity set.
Lemma 2.
For each $R\in\mathcal{R}^{N}$, and each agent $i\in N$, $e_{i}(P)\in T_{i}(P)$.
Proof of Lemma 2.
Let $R\in\mathcal{R}^{N}$, suppose that $TTC(R)$ terminates after Step $K$. Next we prove the lemma by the induction on $\tau=1,\ldots,K$.
Induction basis. For each center $c\in C$, consider the agent in $c$ who has the highest priority according to $\succ_{c}$. By the definition of $\mathcal{T}$, we see that his trading opportunity set equals the set of objects in his center, $O_{c}$. By the definition of TTC, for each $\mathcal{C}\in\mathcal{C}_{1}(R)$, and each $i\in S_{\mathcal{C}}$, we know that (1) $i$ has the highest priority according to $\succ_{c(i)}$; and (2) $e_{i}(R)\in O_{c(i)}$. Thus, for each $i\in S_{\mathcal{C}}$, $e_{i}(R)\in T_{i}(R)=O_{c(i)}$.
Induction hypothesis. Let $L\in\{2,\ldots,K\}$, suppose that for each $\tau<L$, each $\mathcal{C}\in\mathcal{C}_{\tau}(R)$, and each $i\in S_{\mathcal{C}}$, we have $e_{i}(R)\in T_{i}(R)$.
Before we go to the induction step, observe that for $\tau<L$, if an agent $i$ and $e_{i}(R)$ is removed at step $\tau$ under $TTC(R)$, then they are also removed at step $\tau$ under $\mathcal{T}$, according to $R$. This is because under both algorithms, at step $\tau$, $i$ points to his most-preferred remaining object according to $R_{i}$ and $e_{i}(R)$ points to $i$ at step $\tau$. Thus, $i$ and $e_{i}(R)$ are involved in a removed cycle at step $\tau$ under both algorithms as well.
Induction step. Let $\mathcal{C}\in\mathcal{C}_{L}(R)$ and $i\in S_{\mathcal{C}}$. Recall that at the beginning of step $L$ under $TTC(R)$, the remaining objects are $O^{L}$. Hence, $T_{i}(R)\subseteq O^{L}\cap O_{c(i)}$. By the definition of TTC, we know that at step $L$, (1) $e_{i}(R)$ belongs to $c(i$) and is remaining, i.e., $e_{i}(R)\in O^{L}\cap O_{c(i)}$; and (2) agent $i$ has the highest priority among the remaining agents in center $c(i)$ according to $\succ_{c(i)}$. Thus, by the definition of $\mathcal{T}$ and the observation above, we see that all remaining objects in $O^{L}\cap O_{c(i)}$ point to agent $i$ at step $L$ of $\mathcal{T}$, according to $R$. As a result, $e_{i}(R)$ points to $i$ under $\mathcal{T}$. Since $\mathcal{C}$ and $i$ are arbitrary chosen, this argument holds for each agent involved in $\mathcal{C}$. Thus, we conclude that for each $\mathcal{C}\in\mathcal{C}_{L}(R)$, and each involved agent $i\in S_{\mathcal{C}}$, $e_{i}(R)$ to points to $i$ under $\mathcal{T}$. Again, by the observation above, given $R$, we see that all agents involved in $\mathcal{C}$ also form a cycle at step $L$ under $\mathcal{T}$. Therefore, for each $i\in S_{\mathcal{C}}$, $e_{i}(R)\in T_{i}(R)$. ∎
Together with Lemma 2, procedural fairness of $f$ implies that
$\text{for each }R\in\mathcal{R}^{N},\text{ and each }i\in N,f_{i}(R)\mathbin{R_{i}}e_{i}(R).$ (1)
Otherwise $i$ can be better off by consuming $e_{i}(R)\in T_{i}(R)$. Given this fact, for each $R\in\mathcal{R}^{N}$ and each $i\in N$, we say an object is acceptable at $R_{i}$ if $o\mathbin{R_{i}}e_{i}(R)$.
Now, we adapt Sethuraman (2016)’s approach to complete the proof of Theorem 2. For each $R\in\mathcal{R}^{N}$, the size of $R$, $s(R)$, is defined as the total number of acceptable objects across all agents at $R$, i.e., $s(R)=\sum_{i\in N}|\{o\in O\mid o\mathbin{R_{i}}e_{i}(R)\}|$.
[…]
Let $I_{f}(R)=\{i\in N\mid f_{i}(R)\mathbin{P_{i}}TTC_{i}(R)\}$ be the set of agents who prefer $f$ to $TTC$ at $R$. Similarly, $I_{TTC}(R)$ is defined analogously.
Claim 2.
For each $i\in I_{f}(R)\cup I_{TTC}(R)$, agent $i$ has exactly two acceptable objects at $R_{i}$.
Proof of Claim 2.
By contradiction, it is without loss of generality to assume that there exists an agent $i\in I_{f}(R)$ who has more than two acceptable objects at $R_{i}$. Let $R^{\prime}_{i}:f_{i}(R),e_{i}(R),\ldots$ and $R^{\prime}\equiv(R^{\prime}_{i},R_{-i})$.
By strategy-proofness of $f$, $f_{i}(R)=f_{i}(R^{\prime})$; otherwise agent $i$ has an incentive to misreport $R_{i}$ at $f(R^{\prime})$. Since $i\in I_{f}(R)$, $f_{i}(R)\mathbin{P_{i}}TTC_{i}(R)$. Thus, by strategy-proofness of $TTC$, $TTC_{i}(R^{\prime})\neq f_{i}(R)$; otherwise $i$ has an incentive to misreport $R^{\prime}_{i}$ at $TTC(R)$. Overall, $f_{i}(R^{\prime})=f_{i}(R)\neq TTC_{i}(R^{\prime})$. Thus, $R^{\prime}$ is discordant such that $s(R^{\prime})<s(R)$, contradicting the choice of $R$. ∎
By Pycia and Ünver (2017), we know that TTC is Pareto efficient, and hence $I_{TTC}(R)$ is non-empty, otherwise it means $f$ Pareto dominates TTC at $R$.
Let $i\in I_{TTC}(R)$, and let $\mathcal{C}\in\mathcal{C}(R)$ be the cycle involving agent $i$. In other words, $i\in S_{\mathcal{C}}$.
By Claim 2, we know that $R_{i}$ is such that $R_{i}:TTC_{i}(R),e_{i}(R),\ldots$
Since $i\in I_{TTC}(R)$, $\mathcal{C}$ is not executed at $f(R)$. Thus, $f_{i}(R)\neq TTC_{i}(R)$. By (1), $f_{i}(R)\in\{e_{i}(R),TTC_{i}(R)\}$. Therefore, $f_{i}(R)=e_{i}(R)$. This implies that the agent who points to $e_{i}(R)$ in $\mathcal{C}$, say agent $j$, does not receive $e_{i}(R)$ at $f(R)$. By applying similar arguments to agent $j$, we see that $f_{j}(R)=e_{j}(R)$. By repeating this argument for all agents in $S_{\mathcal{C}}$, we conclude that for each agent $k\in S_{\mathcal{C}}$, $f_{k}(R)=e_{k}(R)$. Consequently, $S_{\mathcal{C}}\subseteq I_{TTC}(R)$.
Without of loss of generality, let $S_{\mathcal{C}}=\{i_{1},\ldots,i_{L}\}$. If $L=1$ then $TTC_{i_{1}}(R)=e_{i_{1}}(R)$. Since $f_{i_{1}}(R)\mathbin{R_{i_{1}}}e_{i_{1}}(R)$ we must have $f_{i_{1}}(R)=e_{i_{1}}(R)=TTC_{i_{1}}(R)$, contradicts $S_{\mathcal{C}}\subseteq I_{TTC}(R)$. Thus, $L>1$.
By the definition of TTC, we know that all agents in $S_{\mathcal{C}}$, are form the different centers. Thus, there is a set that consists of $L$ different centers $\{c({i_{1}}),\ldots,c({i_{L}})\}$. For each $i_{\ell}\in S_{\mathcal{C}}$, $TTC_{i_{\ell}}(R)=e_{i_{\ell+1}}(R)\mathbin{P_{i_{\ell}}}e_{i_{\ell}}(R)=f_{i_{\ell}}(R)$ (mod $L$). Also, Lemma 2 implies that $e_{i_{\ell}}(R)\in T_{i_{\ell}}(R)$. Therefore, $f(R)$ is unfairly produced, as agents $\{i_{1},\ldots,i_{L}\}$ can be better off by exchanging from their trading opportunity sets, contradicts procedural fairness of $f$. This completes the proof of Theorem 2.
5 Relation to queuewise rationality
Ekici (2024b) characterizes TTC through strategy-proofness, pair efficiency, and queuewise rationality. Here, we further discuss the relationship between his result and ours.
By Lemma 1, we know that procedural fairness is mathematically stronger than queuewise rationality. However, their economics interpretations are distinct: procedural fairness is a fairness property based on agents’ trading opportunities, queuewise rationality is a voluntary participation condition.
Next, we show that our other fairness properties, internal fairness and external fairness are logically unrelated to queuewise rationality. To illustrate this, we provide three examples below, each satisfying two of the properties but violating the third.
For simplicity, we focus on the case in which $C=\{c,c^{\prime}\}$, $N_{c}=\{1\}$, $O_{c}=\{o_{1}\}$, $N_{c^{\prime}}=\{2,3\}$, $O_{c^{\prime}}=\{o_{2},o_{3}\}$, and $\succ_{c^{\prime}}:2,3$.
Example 2 (Serial-QR mechanism).
We adapt so-called Multiple-Serial-IR mechanisms introduced by Biró et al. (2022).
Serial-QR Algorithm:
Step 0
: Let $\delta=(1,2,3)$ be an order of $N$. Let $R\in\mathcal{R}^{N}$. Let $Y(0)\subseteq X$ be the set of allocations that satisfy queuewise rationality at $R$.
Step $k\geq 1$:
Let $Y_{k}$ be the set of agent $k$’s s allotments that are compatible with some allocation in Y(k1)Y(k-1), i.e., $Y_{k}$ consists of all $y_{k}$ for which there exists an allocation $x\in Y(k-1)$ such that $x_{k}=y_{k}$. Let $y^{*}_{k}$ be agent $k$’s most preferred allotment in $Y_{k}$, i.e., for each $y_{k}\in Y_{k}$, $y^{*}_{k}\mathbin{R_{k}}y_{k}$. Let $Y(k)\subseteq Y(k-1)$ be the set of allocations in Y(k1)Y(k-1) that are compatible with $y^{*}_{k}$, i.e., $Y(k)$ consists of all $x\in Y(k-1)$ with $x_{k}=y^{*}_{k}$.
Termination:
The algorithm terminates after step 3.
Let $\Delta$ be such that for each $R\in\mathcal{R}^{N}$, $\Delta(R)$ is determined by the algorithm above. By the definition of $\Delta$, it satisfies queuewise rationality and internal fairness. To see it violates external fairness, consider $R$ as follows.
$R_{1}:o_{2},o_{1},o_{3};$
$R_{2}:o_{3},o_{2},o_{1};$
$R_{3}:o_{2},o_{1},o_{3}.$
Note that by queuewise rationality, for each $R^{\prime}_{-2}$, $\Delta_{2}(R_{2},R^{\prime}_{-2})=o_{3}$ and hence agent $2$ does not depend on agent $1$. $\Delta(R)=(o_{2},o_{3},o_{1})$ and hence agent $3$ envies agent $1$. Since agents $1$ and $3$ are from different centers, $o_{2}$ belongs to agent $3$’s center, and agent $2$ with higher priority than $3$ does not depend on agent $1$, we see that agent $3$’s envy is externally justified. $\diamond$
Example 3 (Two steps: SD then TTC).
Given $R$, for each center $c$, first run serial dictatorship over $O_{c}$ with respect to $\succ_{c}$: the highest-priority agent in $N_{c}$ picks up his most preferred object in $O_{c}$, the second-highest priority agent picks his most preferred object among the remaining objects in $O_{c}$, and so on. Denote the resulting allocation by $e(R)$. Then, let $e(R)$ be an “artificial” endowment allocation and run the TTC algorithm with respect to $e(R)$, i.e., let each $e_{i}(R)$ points to $i$. This two-step mechanism satisfies queuewise rationality and external fairness, but violates internal fairness. $\diamond$
Example 4 (A variant of serial dictatorships).
Let $SD$ be the serial dictatorship algorithm with $\delta:1,2,3$. That is, agent $1$ picks up his most preferred object in $O$, agent $2$ picks up his his most preferred object among the remaining objects, and the last object is given to agent $3$.
Let $f$ be such that if agents $1$ and $2$ both have the same most preferred object, and that object is in $O_{c^{\prime}}$, then $f(R)=TTC(R)$; otherwise $f(R)=SD(R)$. $f$ satisfies internal fairness and external fairness, but violates queuewise rationality. $\diamond$
5.1 An alternative characterization
Here, adapted from Chen and Feng (2025)’s vacant object lower bounds, we weaken queuewise rationality to the following property, the center lower bound, which imposes that each agent receives an object that is at least as desirable as the least desirable object owned by the center to which he belong.
For each $R\in\mathcal{R}^{N}$, each $c\in C$, and each $i\in N_{C}$, let $\underline{o}_{i}(R)\in O_{c}$ be the worst object in $O_{c}$ at $R_{i}$. That is, for each $o\in O_{c}$, $o\mathbin{R_{i}}\underline{o}_{i}(R)$.
The center lower bound: For each $R\in\mathcal{R}^{N}$ and each $i\in N$, $f_{i}(R)\mathbin{R_{i}}\underline{o}_{i}(R)$.
It is easy to see that the center lower bound is weaker than queuewise rationality. Also, from examples above, we see that the center lower bound is logically independent with internal fairness and external fairness.
Next, from Theorem 1, we characterize TTC by replacing external fairness with the center lower bound.
Theorem 3.
A mechanism $f$ satisfies strategy-proofness, pair efficiency, internal fairness, and the center lower bound, if and only if $f=TTC$.
The proof is similar to Theorem 1, and hence we omit it. Also note that as in Theorem 1, one can weaken internal fairness to weak internal fairness in Theorem 3.
6 Relation to stability
6.1 Core-stability
In this subsection, by adopting the concept of strict core / core-stability, we further explore TTC from the perspective of cooperative game theory. In Shapley-Scarf housing markets, a set of agents $S\subseteq N$ can block an allocation $x$ by reassigning their own endowments among themselves to make all members of $S$ weakly better off and at least one member strictly better off. The definition of blocking can be decomposed into two conditions: one is the weak dominance condition, which says that for a set of agents $S$ to form blocking, all members in $S$ are weakly better off and at least one member is strictly better off; and another one is the self-autarky condition, which states that for $S$ to form a blocking, each member must bring his own endowment.
Still, the first condition can be directly applied to our context. However, the second condition needs to be modified because in our model, an agent may not own a single object, but rather a set of objects associated with the center to which he belongs. Thus, it is natural to require that what agents can bring to form a blocking based on the underlying priority structure. Aligning with our idea of procedural fairness, here we adjust the self-autarky condition by requiring that each member brings one object from his trading opportunity set.
Let $R\in\mathcal{R}^{N}$. At $R$, a set of agents $S\subseteq N$ properly blocks an allocation $x$ via $y$ if
  • (1)
    for each $i\in S$, $y_{i}\mathbin{R_{i}}x_{i}$ and for some $j\in S$, $y_{j}\mathbin{P_{i}}x_{j}$;
  • (2)
    there exists a one to one function $\omega:S\to O$ such that (2-a) for each $i\in S$, $\omega(i)\in T_{i}(R)$; and (2-b) $\{y_{i}\}_{i\in S}=\{\omega(i)\}_{i\in S}$.
Two points are worth noting about our self-autarky condition. First, in the special case where each center contains only one agent, it reduces to the classical self-autarky condition used in the definition of the core. Second, recall that an agent’s trading opportunity set depends on the preferences of agents with higher priority. This restriction implicitly implies that whether an agent can (or cannot) bring an object to form a blocking depends on the views of higher-priority agents. [^18] Aligning with this, Ekici (2024b) modifies the self-autarky condition in a different way: he requires that for each $i\in S$, all agents who are from the same center and have higher priorities than $i$ must also be in $S$.
An allocation is properly core-stable at $R$ if it is not properly blocked by any set of agents. Such an allocation is also referred to as an proper core allocation. Note that any proper core allocation is fairly produced, but not vice versa. The proper core, $PC(R)$ is the set that contains all proper core allocations at $R$.
We next show that the proper core is single-valued: Particularly, it is induced by TTC.
Theorem 4.
For each $R\in\mathcal{R}^{N}$, $PC(R)=\{TTC(R)\}$.
6.1.1 Proof of Theorem 4
First, we provide one useful result about agents’ trading opportunity sets. For each $R\in\mathcal{R}^{N}$, let $I_{\tau}(R)\subseteq N$ and $O_{\tau}(R)\subseteq O$ be the set of agents and the set of objects that are removed at the end of step $\tau$ under $TTC(R)$, respectively.
Claim 3.
For each $t$, and each $j\in I_{t}$, $T_{j}(R)\cap\bigcup_{\tau=1}^{t-1}O_{\tau}(R)=\emptyset$.
The proof immediately follows the definition of $\mathcal{T}$ and hence we omit it. The rest of proof is similar to the proof in Roth and Postlewaite (1977), who demonstrate that the TTC always induces the unique core allocation in Shapley-Scarf housing markets.
We first show that TTC induces a proper core allocation. Suppose that TTC does not induce a proper core allocation. Then, there are $R\in\mathcal{R}^{N}$ and $S\subseteq N$ such that $S$ properly blocks $TTC(R)$ via an allocation $y\neq TTC(R)$. Let $x\equiv TTC(R)$. Let $\mathcal{C}\in\mathcal{C}_{1}(R)$ be a trading cycle formed at the first step under $TTC(R)$, and recall that $S_{\mathcal{C}}\subseteq N$ is the set of agents who are involved in $\mathcal{C}$. There are two cases. First, if an agent $i\in S_{\mathcal{C}}$ is also in $S$, then we have $y_{i}=x_{i}$ as $x_{i}$ is his most preferred object. Then, by our self-autarky condition (2), there is an agent $\widehat{i}\in S$ such that $x_{i}\in T_{\widehat{i}}(R)$. It implies that $\widehat{i}\in S_{\mathcal{C}}$; otherwise $i$ cannot receive $x_{i}$ via TTC. Generally, we find that for each object involved in $\mathcal{C}$ there is a unique agent whose trading opportunity set contains that object, namely, the agent in $S_{\mathcal{C}}$ to whom the object points. Together with our self-autarky condition (2), it implies that agent is also in $S$. Consequently, we conclude that $S_{\mathcal{C}}\subseteq S$, and all agents in $S_{\mathcal{C}}$ receive their TTC allocation at $y$ as well. This implies that no agent in $S$ can be strictly better off by receiving an object from $\{x_{i}\}_{i\in S_{\mathcal{C}}}$. Second, if no agent in $S_{\mathcal{C}}$ is in $S$, then by Claim 3, we see that for each agent in $S$, his trading opportunity set does not contain any object from $\{x_{i}\}_{i\in S_{\mathcal{C}}}$. In any case, we conclude that no agent in $S$ can be strictly better off by receiving an object from an object that is removed from the first step of TTC. Applying the argument to the second step and so on, we obtain that there is no agent $j\in S$ such that $y_{j}\mathbin{P_{j}}x_{j}$, leading to a contradiction.
The converse can be proved by adopting Lemma 1 in Roth and Postlewaite (1977). Let $R\in\mathcal{R}^{N}$, $x\equiv TTC(R)$, and $\mathcal{C}_{1},\ldots,\mathcal{C}_{K}$ be corresponding trading cycles. Without loss of generality, assume that for any $t,\tau=1,\ldots,K$, $t<\tau$ means that $\mathcal{C}_{t}$ is executed no later than $\mathcal{C}_{\tau}$. We also have corresponding coalitions $S_{\mathcal{C}_{1}},\ldots,S_{\mathcal{C}_{K}}$. Now, consider an arbitrary allocation $y\in X$ with $y\neq x$. Let $K^{*}\in\{1\ldots,K\}$ be the first integer where there exists an agent $j\in S_{\mathcal{C}_{K^{*}}}$ such that $x_{j}\neq y_{j}$ Then, by construction, coalition $S=\cup_{\kappa=1}^{K^{*}}S_{\mathcal{C}_{\kappa}}$ properly blocks $y$ via $x$. To see it, we check our two conditions one by one. By the selection of $K^{*}$, for each $i\in S\setminus S_{\mathcal{C}_{K^{*}}}$, $x_{i}=y_{i}$; by the definition of TTC, for each $i\in S_{\mathcal{C}_{K^{*}}}$ and each $o\in O\setminus\{x_{i}\}_{i\in S\setminus S_{\mathcal{C}_{K^{*}}}}$, $x_{i}\mathbin{R_{i}}o$. Thus, for each $i\in S_{C_{K^{*}}}$, $x_{i}\mathbin{R_{i}}y_{i}$, and hence $x_{j}\mathbin{P_{j}}y_{j}$, as $x_{j}\neq y_{j}\in O\setminus\{x_{i}\}_{i\in S\setminus S_{\mathcal{C}_{K^{*}}}}$. That is, condition (1) is satisfied. By the definition of TTC, we see $\{x_{i}\}_{i\in S}$ satisfies condition (2).
6.2 Pairwise stability
For priority-based allocation problems, e.g., school choice problems, a standard stability notion is pairwise stability (Gale and Shapley, 1962): An allocation $x$ is pairwisely stable at $R$ if there is no agent-object pair $(i,o)$ such that $o\mathbin{P_{i}}x_{i}$ and $i\succ_{o}j$, where $x_{j}=o$. Clearly, this notion is not suitable here, as in our model, objects have no priorities over all agents. Even if we adopt each center $c$’s priority for objects in $O_{c}$, it would still be problematic because $\succ_{c}$ is incomplete. One way to address this issue is to expand the priorities to be complete. A natural expansion is as follows: For each $c\in C$, each $o\in O_{c}$, each $i\in N_{c}$, and each $j\not\in N_{c}$, $i$ has higher expanded priority at $o$ than jj. [^19] Karakaya et al. (2019) propose an extension of TTC based on such expanded priorities, and Combe (2023) shows that this TTC extension is “minimal envy” among the set of individually rational, strategy-proof, and Pareto efficient mechanisms. However, this assumption will lead to an incompatibility with pair efficiency. To see it, let us revisit Example 1. Suppose that at $a$ and $b$, agent $1$ has higher priority than $3$, as $a,b\in O_{c},1\in N_{c}$, and $3\not\in N_{c}$. Similarly, $3$ has higher priority than $1$ and $2$ at $d$. Suppose that $2$ receives $d$: in this case, if $3$ receives $b$, then $3$ would block this allocation as $3$ has higher priority than $2$ at $d$; however, if $3$ receives $a$ then it means $1$ receives $b$. So, $1$ would block this allocation as $1$ has higher priority than $3$ at $a$. Thus, $2$ does not receive $d$. Since $2$ also has priority than $1$ at $a$, we conclude that it must be the case that $2$ receives $a$. Subsequently, $1$ receives $b$ and $3$ receives $d$, i.e., the allocation $y$. However, this allocation is not pair efficient: agents $2$ and $3$ can be better off by switching their allotments.
Therefore, finding alternative ways to expand priorities to be complete and defining criteria to select “fair” expansions constitutes an interesting open problem, though one that is beyond the scope of this paper.
7 Conclusion
In this paper, we investigate Ekici (2024b)’s multi-center allocation problems. By proposing different fairness properties, we provide several TTC characterizations. Our analysis sheds light on Ekici (2024b) to show that TTC performs surprisingly well according to different objectives, e.g., efficiency, fairness, and incentives. We finish this paper with some final remarks related to issues left aside during the presentation of the main results.
7.1 Length constraints
The primary application target of this paper is cross-program cooperation in organ exchange programs, e.g., kidney and liver exchanges. One might wonder if TTC is not applicable for kidney exchanges, as it allows trading cycles without length constraints (of course, in our model, each cycle contains at most $|C|$ agents). [^20] If $|C|$ is sufficiently small, we may not even meet length constraints. Note that in India, cyclic exchanges involving up to ten pairs have been successfully conducted (Kute et al., 2021). In practice, cycle length constraints are imposed due to the scarcity of medical resources: a two-way exchange requires the simultaneous availability of four operating rooms and associated personnel, while a three-way exchange requires six, and so on. However, in our model, the presence of multiple centers allows for integrating medical resources from different centers, making it feasible to conduct exchanges with longer cycles. [^21] There are also studies that examine cooperation across different kidney exchange programs without considering length constraints; see, e.g., Benedek et al. (2024).
7.2 Balancedness
Another restriction in our model is balancedness: For each center $c$, the inflow of agents coming from other centers to receive objects in $O_{c}$ is equal to the outflow of agents in $N_{c}$ who leave $c$ to receive objects in other centers. This restriction aligns with real-world scenarios in kidney exchanges: As in the classical kidney exchange model, assuming $|N_{c}|=|O_{C}|$ is based on the observation that, in most cases, each patient comes with a living kidney donor, often a family member. [^22] For organ exchange environments with imbalance and capacity constraints, Ekici (2024b) proposes a solution by introducing fictitious agents and objects; the same construction can be applied here as well. While this restriction seem to limit the applicability of this paper to other use cases, we argue this point by the following reason: Balancedness is an inviolable constraint in many practical allocation problems, such as shift reallocation and exchange programs (Dur and Ünver, 2019; Yu and Zhang, 2020; Manjunath and Westkamp, 2021; Kamada and Kojima, 2025). As pointed out by Dur and Ünver (2019), in such settings, every position and every worker must be matched, and hence the balancedness restriction is imposed. Thus, balancedness does not reduce the applicability of this paper. Of course, considering extensions involving imbalance or varying object capacities is an interesting open problem. However, such issues are beyond the scope of this paper, and we leave them for future research.
Appendix A Proofs
A.1 Proof of Lemma 1
Let $f$ be procedurally fair. Let $R\in\mathcal{R}^{N}$ and $x\equiv f(R)$. Let $i\in N$, $c\equiv c(i)$. Suppose that $i$ has the $k$-highest priority at $c$, i.e., $r_{i}=k$. We see that at most k1k-1 objects in $O_{c}$ are not in $T_{i}(R)$. Consider the worst case that all the k1k-1 number of objects are more preferred than $o_{i}(R)$. In this case, $o_{i}(R)\in T_{i}(R)$. Therefore, $x_{i}\mathbin{R_{i}}o_{i}(R)$ holds, because at least agent $i$ can directly consume $o_{i}(R)$ from $T_{i}(R)$.
Next, we use an example to show that a queuewise rational allocation may not be fairly produced.
Let $C=\{c_{1},c_{2}\}$, $N_{c_{1}}=\{1,2\},N_{c_{2}}=\{3\}$, $O_{c_{1}}=\{o_{1},o_{2}\},O_{c_{2}}=\{o_{3}\}$, and $\succ_{c_{1}}:1,2$.
Let $R$ be such that
$R_{1}:o_{3},o_{1},o_{2};$
$R_{2}:o_{3},o_{2},o_{1};$
$R_{3}:o_{1},o_{2},o_{3}.$
Let $y\equiv(o_{1},o_{2},o_{3})$. We see that $y$ is queuewise rational at $R$. Note that $T_{1}(y)=\{o_{1},o_{2}\}$ and $T_{3}(y)=\{o_{3}\}$. We find that agents $1$ and $3$ can be better off by trading $(o_{1},o_{3})\in T_{1}(y)\times T_{3}(y)$. Thus, $x$ is unfairly produced at $R$.
A.2 Proof of Proposition 1
Recall that TTC is pair efficient and strategy-proof as it is a sub-class of trading cycle mechanisms (Pycia and Ünver, 2017).
Internal fairness. Suppose that TTC violates internal fairness. Then, there is a profile $R$, a center $c\in C$, and a pair of agents $i,j\in N_{C}$ such that $i\succ_{c}j$, and $i$ envies $j$ at $TTC(R)$. If $TTC_{j}(R)\in O_{c}$, then by the definition of TTC, we know that $i$ can receive $TTC_{j}(R)$ by pointing to $TTC_{j}(R)$ as $i\succ_{c}j$. If $TTC_{j}(R)\not\in O_{c}$. Let $R^{\prime}_{i}:TTC_{j}(R),\ldots$ We see that $TTC_{i}(R^{\prime}_{i},R_{-i})=TTC_{j}(R)$ by the definition of TTC, as $i$ moves earlier than $j$. This implies that TTC is not strategy-proof, a contradiction.
External fairness. By the definition of TTC, we know that each agent in a cycle depends on each of the others in that cycle (or agents in the previous cycles). Also we know that, if $i$ envies $j$, then $j$ receives his allotment, denoted by $o$, at an earlier step than $i$. Let $\mathcal{C}$ be the cycle that includes $j$. So, $i\in N_{c(i)}$ and $j\not\in N_{c(i)}$, and recall that $o\in O_{c(i)}$. In $\mathcal{C}$, $o$ is pointing to an agent $k\in N_{c(i)}$ who is higher ranked than $i$ (i.e., $k\succ_{c(i)}i$) and who depends on $j$. Therefore, this envy is not externally justified.
Procedural fairness. Finally, we inductively show that TTC is procedurally fair. Let $R\in\mathcal{R}^{N}$. Suppose that $TTC(R)$ stops at step $K$.
Induction basis. Let $\mathcal{C}\in\mathcal{C}_{1}(R)$ and $i\in S_{\mathcal{C}}\subseteq I_{1}$. By the definition of TTC, we know that $x_{i}$ is $i$’s most preferred object. Thus, we conclude no agent in $I_{1}$ can be better off either by directly consuming an object from his trading opportunity set or by trading with other agents.
Induction hypothesis. Let $\kappa\leq K$. Suppose that for $\tau=2,\ldots,\kappa-1$, no agent in $I_{\tau}$ can be better off either by directly consuming an object from his trading opportunity set or by trading with other agents.
Induction step. We show that no agent in $I_{\kappa}$ can be better off either by directly consuming an object from his trading opportunity set or by trading with other agents. Let $O^{\kappa}$ be the set of remaining objects at the beginning of step $\kappa$ under $TTC(R)$. Let $i\in I_{\kappa}$.
Case (a). Direct consumption. By the definition of TTC, $TTC_{i}(R)$ is $i$’s most preferred object in $O^{\kappa}$. By Claim 3, $T_{i}(R)\subseteq{O}^{\kappa}$. Thus, $i$ cannot be better off by directly consuming an object from his trading opportunity set $T_{i}(R)$.
Case (b). Exchange. By contradiction, suppose that $i$ can be better off by trading with other agents. Thus, there is a list of centers $c_{1},\ldots,c_{L}\in C$, a list of agents $(i_{1},\ldots,i_{L})\in\prod_{\ell=1}^{L}N_{c_{\ell}}$, and a list of objects $(o_{1},\ldots,o_{L})\in\prod_{\ell=1}^{L}T_{i_{\ell}}(R)$ such that for each $\ell\in\{i_{1},\ldots,i_{L}\}$, $o_{\ell+1}\mathbin{P_{\ell}}TTC_{\ell}(R)$ (mod $L$). Without of loss of generality, let $i_{1}=i$. By Case (a) we know that $L>1$. Thus, $i_{2}\neq i$. Now consider $i_{2}$. By the induction hypothesis, no agent in $\cup_{\tau=1}^{\kappa-1}I_{\tau}$ can be better off by trading with $i$. Thus, $i_{2}$ is remaining at step $\kappa$, i.e., $i_{2}\not\in\cup_{\tau=1}^{\kappa-1}I_{\tau}$. Then, by Claim 3, we have $T_{i_{2}}(R)\subseteq{O}^{\kappa}$. Also, recall that for each $o\in O^{L}$, $TTC_{i}(R)\mathbin{R_{i}}o$. Therefore, $i$ cannot be better off by trading with $i_{2}$.
Appendix B Independence examples
We establish the independence of properties through the following examples. We label examples by the property that is not satisfied.
Theorem 1
Example 5 (Staregy-proofness).
Let $C=\{c_{1},c_{2},c_{3}\}$, $N_{c_{1}}=\{1\}$, $O_{c_{1}}=\{o_{1}\}$, $N_{c_{2}}=\{2\}$, $O_{c_{2}}=\{o_{2}\}$, $N_{c_{3}}=\{3,4\}$, $O_{c_{3}}=\{o_{3},o_{4}\}$, and $\succ_{c_{3}}:3,4$.
Consider a “circular preference profile” as the one below.
$\widetilde{R}_{1}:o_{2},o_{3},o_{1},o_{4};$
$\widetilde{R}_{2}:o_{3},o_{1},o_{2},o_{4};$
$\widetilde{R}_{3}:o_{1},o_{2},o_{3},o_{4}.$
$\widetilde{R}_{4}:\ldots$
Note that $x\equiv(o_{3},o_{1},o_{2},o_{4})$ is pair efficient at $\widetilde{R}$. However, TTC(R~)TTC($\widetilde${R}) outputs the allocation $(o_{2},o_{3},o_{1},o_{4})$, which makes agents $\{1,2,3\}$ strictly better off.
Let $f$ be such that $f(\widetilde{R})=x$ and, for each $R\neq\widetilde{R}$, let $f(R)=TTC(R)$. $f$ is pair efficient, internally fair, externally fair, but is not strategy-proof. $\diamond$
Example 6 (Pair efficiency).
Let $C=\{c_{1},c_{2}\}$, $N_{c_{1}}=\{1\}$, $O_{c_{1}}=\{o_{1}\}$, $N_{c_{2}}=\{2,3\}$, $O_{c_{2}}=\{o_{2},o_{3}\}$, and $\succ_{c_{2}}:,2,3$.
Let $f$ be such that for each $R\in\mathcal{R}^{N}$, $f_{1}(R)=o_{1}$ and $f_{2}(R)\mathbin{R_{2}}f_{3}(R)$. Essentially, it means that agent $1$ always receives his own, and we run a serial dictatorship algorithm with $2,3$. We see that $f$ violates pair efficiency and satisfies other three properties. $\diamond$
Example 7 (Internal fairness).
Let $|C|=1$. Thus, external fairness is trivially satisfied. Run the serial dictatorship algorithm that does not respect the center’s priority $\succ_{c}$. It is easy to see that it is strategy-proof and pair efficient but is not internally fair. $\diamond$
Example 8 (External fairness).
Let $C=\{c_{1},c_{2}\}$, $N_{c_{1}}=\{1\}$, $O_{c_{1}}=\{o_{1}\}$, $N_{c_{2}}=\{2,3\}$, $O_{c_{2}}=\{o_{2},o_{3}\}$, and $\succ_{c_{2}}:,2,3$.
Let SD be the serial dictatorship with $\pi:2,1,3$. It is easy to see that SD is strategy-proof, pair efficient, and internally fair but is not externally fair. $\diamond$
Theorem 2 and Procedural fairness
We first show that strategy-proofness and procedural fairness are independent.
Staregy-proofness: Consider a modification of Example 5 as follows.
Let $C=\{c_{1},c_{2},c_{3}\}$, $N_{c_{1}}=\{1\}$, $O_{c_{1}}=\{o_{1}\}$, $N_{c_{2}}=\{2\}$, $O_{c_{2}}=\{o_{2}\}$, $N_{c_{3}}=\{3,4\}$, $O_{c_{3}}=\{o_{3},o_{4}\}$, and $\succ_{c_{3}}:3,4$.
Let $\widetilde{R}$ be such that
$\widetilde{R}_{1}:o_{2},o_{3},o_{1},o_{4};$
$\widetilde{R}_{2}:o_{1},o_{3},o_{2},o_{4};$
$\widetilde{R}_{3}:o_{1},o_{2},o_{3},o_{4}.$
$\widetilde{R}_{4}:\ldots$
Let $f$ be such that $f(\widetilde{R})=(o_{2},o_{3},o_{1},o_{4})$ and, for each $R\neq\widetilde{R}$, $f(R)=TTC(R)$. $f$ is procedurally fair but is not strategy-proof. $\diamond$
Procedural fairness: A constant mechanism that always selects the same allocation is strategy-proof but is not procedurally fair. $\diamond$
Next, we show that procedural fairness is independent with other mentioned properties. By Theorem 2 and independence examples in Theorem 1, one can find that pair efficiency / internal fairness / external fairness cannot imply procedural fairness solely. Thus, we only show that procedural fairness does not imply any of the other properties.
Pair efficiency: Let $C=\{c_{1},c_{2},c_{3}\}$, $N_{c_{1}}=\{1,2\},N_{c_{2}}=\{3\},N_{c_{3}}=\{4\}$, $O_{c_{1}}=\{o_{1},o_{2}\},O_{c_{2}}=\{o_{3}\},O_{c_{3}}=\{o_{4}\}$, and $\succ_{c_{1}}:1,2$.
Let $R$ be such that
$R_{1}:o_{3},o_{4},\ldots;$
$R_{2}:o_{4},o_{3},\ldots;$
$R_{3}:o_{1},o_{2},\ldots;$
$R_{4}:o_{2},o_{1},\ldots.$
Consider $x\equiv(o_{3},o_{4},o_{2},o_{1})$. Note that $T_{1}(R)=\{o_{1},o_{2}\},T_{3}(R)=\{o_{3}\},T_{4}(R)=\{o_{4}\}$, and $T_{2}(R)=\{o_{2}\}$. One can see that agents $1$ and $2$ cannot be better off because they receives their most preferred objects. Also, agents $3$ and $4$ prefer any object in $\{o_{1},o_{2}\}$ over any object in $\{o_{3},o_{4}\}$. Thus, $x$ is fairly produced. However, $x$ is not pair efficient as agents $3$ and $4$ can be better off by switching their allotments. $\diamond$
Internal fairness: Let $C=\{c_{1},c_{2}\}$, $N_{c_{1}}=\{1,2\},N_{c_{2}}=\{3\}$, $O_{c_{1}}=\{o_{1},o_{2}\},O_{c_{2}}=\{o_{3}\}$, and $\succ_{c_{1}}:1,2$.
Let $R$ be such that
$R_{1}:o_{3},o_{2},o_{1};$
$R_{2}:o_{3},o_{2},o_{1};$
$R_{3}:o_{1},o_{2},o_{3}.$
Consider $\equiv(o_{2},o_{3},o_{1}). xx$ is not internally fair as $1$ envies $2$ and $1\succ_{c_{1}}2$. However, $x$ is fairly produced. To see it, first note that $T_{1}(x)=\{o_{1},o_{2}\},T_{3}=\{o_{3}\}$, and $T_{2}(x)$ is either $\{o_{1}\}$ or $\{o_{2}\}$. So, while agent $1$ prefers $o_{2}$ to $o_{1}$, $1$ cannot be better off solely via consuming another object from $T_{1}(x)$. Also note that only agent $1$ does not receive his most preferred object. It means that agent $1$ cannot be better off by trading with other agents. $\diamond$
External fairness: Consider a modification of Serial-QR mechanism in Example 2, where $Y(0)$ is adjusted to include all fairly produced allocations. This mechanism is procedurally fair but is not externally fair. $\diamond$
Theorem 3
Examples 4, 5, 6, and 7 show the independence of Theorem 3.